geerky42 (geerky42):

Show that \(\sin^{-1}(x) + \cos^{-1}(x) = \dfrac{\pi}{2}\) for \(|x| \leq 1\)

geerky42 (geerky42):

What identifies would be great for me to use? I just need me some leads...

geerky42 (geerky42):

derivative? it would get me to 0 = 0. ok so I know arcsin(x) + arccos(x) should be equal to constant, but why pi/2 ?

geerky42 (geerky42):

@Loser66

OpenStudy (loser66):

I don't know, but why not pi/2? it's a number, still

geerky42 (geerky42):

we can't just make up what it equal to, that's not really math lol...

OpenStudy (aravindg):



OpenStudy (aravindg):

x+y=pi/2 y=pi/2-x Put sin x=k sin y=l sin^-1 k+sin^-1 l=x+pi/2-x=pi/2

OpenStudy (aravindg):

Is that explanation convincing? :)

geerky42 (geerky42):

hmm yeah. good enough lol. thanks! I was just curious.

OpenStudy (aravindg):

yw :)